### v1

# code[Code][List of Codes]

Abstract
Fine-grained and contact-rich manipulation remain challenging for robots, largely due to the underutilization of tactile feedback. To address this, we introduce TouchGuide, a novel cross-policy visuo-tactile fusion paradigm that fuses modalities within a low-dimensional action space. Specifically, TouchGuide operates in two stages to guide a pre-trained diffusion or flow-matching visuomotor policy at inference time. First, the policy produces a coarse, visually-plausible action using only visual inputs during early sampling. Second, a task-specific Contact Physical Model (CPM) provides tactile guidance to steer and refine the action, ensuring it aligns with realistic physical contact conditions. Trained through contrastive learning on limited expert demonstrations, the CPM provides a tactile-informed feasibility score to steer the sampling process toward refined actions that satisfy physical contact constraints. Furthermore, to facilitate TouchGuide training with high-quality and cost-effective data, we introduce TacUMI, a data collection system. TacUMI achieves a favorable trade-off between precision and affordability; by leveraging rigid fingertips, it obtains direct tactile feedback, thereby enabling the collection of reliable tactile data. Extensive experiments on five challenging contact-rich tasks, such as shoe lacing and chip handover, show that TouchGuide consistently and significantly outperforms state-of-the-art visuo-tactile policies.
Figure 1: TacUMI is a low-cost yet high-precision handheld data collection system that provides direct tactile feedback through a rigid mechanical coupling. TouchGuide is a multi-modal fusion paradigm that steers a visuomotor policy via touch guidance during denoising or flow matching, producing actions that better adhere to contact physics without retraining the base policy.
I Introduction
Imitation learning [75, 69, 14, 44, 56] has enabled robots to perform complex manipulation tasks [47], such as folding clothes and preparing coffee [31, 45, 5, 3]. However, many intuitive fine-grained and contact-rich skills (e.g., shoe lacing or handing over fragile chips) remain challenging. These tasks demand a policy that effectively integration of visual and tactile information: visual information provides global scene context and task understanding, while tactile feedback supplies localized, contact-specific information essential for fine-grained control. This motivates a central question: How should visual and tactile information be fused to improve policy planning in fine-grained and contact-rich manipulation?
Existing visuo-tactile policy models generally follow one of two dominant fusion strategies: feature-level [65, 57, 13, 4] or policy-level [9, 8] fusion. While policy-level methods can mitigate the modality-dominance issue in feature-level concatenation caused by sparse features, they typically require training multiple single-modality policies, which makes it difficult to capture the cross-modal correlations that are critical for tasks demanding tight multimodal synergy.
To address this fundamental limitation, we introduce TouchGuide, a novel cross-policy visuo-tactile fusion paradigm. Departing from conventional computer vision approaches that focus on “fusing multi-modal data per se”, we rethink the problem from a robotic control perspective. Our core idea is to extract task-relevant semantics from each modality and perform fusion within the low-dimensional action space in a way that can benefits robotic manipulation. Concretely, TouchGuide steers a pre-trained visuomotor policy based on diffusion or flow matching (e.g., the action expert in $\pi_{0.5}$) during inference. The process begins by generating a visually feasible, coarse action from visual observation alone in the early sampling phase. Then, in the later phase, a task-specific Contact Physical Model (CPM) injects tactile guidance, refining the action to better satisfy real-world physical contact constraints. The CPM is trained via contrastive learning on limited expert demonstrations to directly output a feasibility score. This score assesses whether the action comply with the contact physical constraints implied by the current tactile observations in task-specific environment, thereby generating a more robust steering signal. Notably, CPM can be applied across multiple policies, including VLAs [32, 5, 50, 58], diffusion policies [14, 69, 56, 44], and flow-matching policies [34, 71].
Moreover, in fine-grained and contact-rich manipulation, policy execution performance depends heavily on the quality of expert demonstrations. Teleoperation, due to high latency and indirect haptic feedback, hinders intuitive control. While handheld devices such as UMI [15] provide a platform-decoupled solution for data collection, existing systems relying on SLAM-based localization or motion capture (MoCap) cannot simultaneously achieve both high accuracy and low cost. To this end, we introduce TacUMI, designed according to the following principles: (1) Cost-effective (\$720 without tactile sensors) and lightweight (540g), while providing the high-precision data required for fine-grained manipulation. (2) Direct tactile feedback enabled by rigid fingertips. (3) A unified interface for synchronized acquisition of high-resolution visual and tactile signals.
Our main contributions are threefold:
  • We propose TouchGuide, a novel cross-policy fusion paradigm for fine-grained and contact-rich manipulation.
  • We develop TacUMI, a cost-effective yet high-precision data collection system with direct tactile feedback.
  • Experiments on five challenging fine-grained and contact-rich tasks show that our TouchGuide can outperform other SOTA visuo-tactile methods and generalize across policies and tactile modalities effectively.
II Related Work
Figure 2: Overview of TacUMI data collection system. TacUMI (Collection-side) uses a Vive tracker for localization to obtain accurate end-effector poses, while the operator receives direct tactile feedback (Left). During policy inference, we use an execution-side device that is structurally identical to the collection-side TacUMI, coupled to different robot arms via an adapter.
II-A Hardware and Data Collection System
II-A1 Tactile Sensors
Optical tactile sensors [68, 53, 37, 49, 41, 40, 36, 62] have advanced rapidly in recent years and have been widely adopted, owing to their ability to provide rich information (i.e., force and torque can be indirectly inferred from tactile images [74, 73]). Considering cost-effectiveness and ease of deployment, we selected the Xense tactile sensor [62]. The sensor offers software force estimation interfaces [61], making it possible to evaluate our system algorithms across multiple tactile modalities.
II-A2 Data Collection System
Teleoperation is a common choice for data collection, from ALOHA-style leader–follower systems [75, 46, 1, 10, 60, 25, 6] to peripheral-driven control such as 3D mice [19], VR controllers [65, 33], and hand tracking [48]. However, contact-rich and precise manipulation remains challenging due to missing tactile feedback; existing solutions mainly provide indirect haptic cues via VR visualization [65, 35, 20].
Some UMI-like [15, 76] handheld gripper data collection devices offer a data collection paradigm that operates independently of the physical robot [79, 43, 76, 63]. This type of handheld data-collection gripper, which operates without the physical robot, can provide haptic feedback directly, and many tactile data collection systems use this paradigm [64, 42, 79, 59, 67, 28, 39, 16]. ViTaMIn [42] and Touch in the Wild [79] adopt SLAM-based localization; however, the resulting accuracy is insufficient for many precise tasks. Some works (e.g., exUMI [64], ViTaMIn-B [39]) attach a VR controller to the gripper for localization, but this also introduces a notable drawback: it requires wearing a VR headset, which can impose a substantial burden during long-duration data collection. FreeTacMan [59] and FARM [28] use motion-capture systems, achieving high accuracy while remaining lightweight. However, they impose stricter requirements on the capture environment, and the high cost of motion-capture equipment makes large-scale data collection difficult.
In contrast, our TacUMI uses Vive Trackers [54] for localization, reducing cost as much as possible while still meeting the accuracy requirements of precise manipulation tasks, thereby enabling large-scale data collection. Meanwhile, we minimize the linkage structure between the fingers and the gripper to obtain more direct haptic feedback, improving the quality of the tactile data.
II-B Visuo-tactile Policies
Visuo-tactile policies are crucial for precise manipulation, and prior works [29, 65, 70, 13, 27, 30, 4, 52, 21, 38, 12, 26, 66, 17, 9] have continuously explored this direction. Most of these policies (e.g., 3D-ViTac [29], VTLA [70], FreeTacMan [59]) rely on simple feature-level concatenation, which can cause the policy to overemphasize a single modality [9]. Policy-level composition methods such as PolicyConsensus [9] and stage-wise fusion approaches like RDP [65] alleviate this issue to some extent, but they do not perform multi-modal fusion in the action space. In contrast, our method, TouchGuide, introduces Contact Physical Model (CPM) to efficiently fuse modalities in the action space, steering the policy distribution toward the real distribution (See Sec. IV-C).
II-C Inference-time Steering
Inspired by classifier guidance [18], some recent works [22, 51, 55, 72] have begun to adopt inference-time steering for robotic manipulation, which is an effective way to guide actions toward a target distribution during inference. DynaGuide [22] and LPB [51] guide visuomotor policies via a dynamics model; however, this paradigm is ineffective for fine-grained manipulation, because future observations can become entirely different due to small errors in the action. In contrast, TouchGuide focuses on the current contact physics and can be integrated with a wide range of visuomotor policies (e.g., DP [14] and $\pi_{0.5}$ [32]), yielding clear improvements over dynamics-based approaches (Results are presented in Sec. V).
III Data Collection System: TacUMI
III-A System Overview
As shown in Fig. 2, TacUMI is a UMI-like [15] handheld gripper that collects tactile data without relying on proprioception. Its development follows three key characteristics:
Accuracy and Low-cost. TacUMI is localized using a Vive tracker with two Lighthouse base stations, providing reliable and accurate pose estimates. In contrast, UMI [15] relies on SLAM-based tracking; for fine-grained tactile manipulation tasks which requires high precision, its accuracy is insufficient and tracking can be easily lost in feature-sparse scenes. Notably, a setup with one Vive tracker and two Lighthouse base stations costs only about \$720 (without tactile sensors), maintaining a low cost while providing accurate localization. A detailed comparison of data collection quality across different methods will be presented in the experiments (See Sec. V-D).
Direct Tactile Feedback. Given that collecting tactile data for fine-grained manipulation critically depends on real-time tactile feedback, teleoperation presents a fundamental challenge by introducing a delay that impedes the operator’s ability to perceive and respond to touch signals promptly. Although some teleoperation methods, such as TactAR [65], provide visualized tactile feedback, this relies on human learning. After becoming accustomed to the interface, operators often ignore the visual tactile cues and instead act based on experience [65]. In contrast, this gripper-finger design enables direct tactile feedback, facilitating the collection of high-quality data.
Figure 3: Overview of TouchGuide framework. (a) The architecture of the task-specific Contact Physical Model (CPM). (b) During inference, the CPM serves as an external model that steers the base policy’s action generation within the sampling process using a feasibility score. (c) In action space, TouchGuide can be viewed as a form of contact-physics steering that steers the policy distribution toward the real distribution.
Usability and Synchronization. The entire gripper adopts a unified hardware-software interface. By connecting it to a computer via a single aviation cable, we can stream observations, including tactile and visual signals as well as proprioceptive states. In addition, our data collection is compatible with the LeRobot [7] format and uses a unified software interface to acquire all modalities, thereby achieving effective software-level synchronization. Moreover, TacUMI captures high-resolution visual $\mathbf{V}_{t}\in\mathbb{R}^{640\times 480\times 3}$ and tactile $\mathbf{T}_{t}\in\mathbb{R}^{200\times 350\times 3}$ images at 30Hz, which suffices for the vast majority of fine-grained manipulation tasks. Meanwhile, we integrate an RK3576 CPU inside TacUMI, which enables on-board computation of force while capturing tactile images. Due to the modality-specific nature of tactile sensing (i.e., different tactile sensors cannot share the same tactile images for training), this capability is crucial for tactile data collection.
More details on hardware selection, design choices, and system parameters are provided in the Appendix -D.
III-B Data Processing for Training
We define $\{W\}$ as the world frame provided by Vive-tracker Lighthouse, $\{T\}$ as the Vive-tracker frame, $\{E\}$ as the end-effector (ee) frame and $\{B\}$ as the robot base frame. Before data collection, we align the robot base frame with the world frame, i.e., the transformation matrix $^{W}T_{B}\in SE(3)$ should be obtained, and $^{T}T_{E}$ can be computed from the 3D model file. The data sequence collected by the Vive tracker after the coordinate transformation is $\{^{B}T_{E}^{(0)}\dots{}^{B}T_{E}^{(i)}\}$, i.e., the end-effector pose in the robot base frame. Similar to FastUMI [76], the data sequence can be expressed as $\{\mathbf{p}^{(i)}_{\text{ee}},\mathbf{R}^{(i)}_{\text{ee}}\}$, where $\mathbf{p}^{(i)}_{\text{ee}}$ and $\mathbf{R}^{(i)}_{\text{ee}}$ are the end-effector position and orientation in the robot base frame, respectively.
For policy training, we use relative ee pose $\{\mathbf{p}_{\text{rel}}^{(i)},\mathbf{R}_{\text{rel}}^{(i)}\}$ as the dataset format, which is considered more effective than using absolute ee pose [15]. This can be expressed as follows:
$\mathbf{p}_{\text{rel}}^{(i)}$ $={\mathbf{R}_{\text{ee}}^{(0)\top}}(\mathbf{p}_{\text{ee}}^{(i)}-\mathbf{p}_{\text{ee}}^{(0)})$ (1)
$\mathbf{R}_{\text{rel}}^{(i)}$ $={\mathbf{R}_{\text{ee}}^{(0)\top}}\mathbf{R}_{\text{ee}}^{(i)}$ (2)
where $\{\mathbf{p}_{\text{ee}}^{(0)},\mathbf{R}_{\text{ee}}^{(0)}\}$ denotes the last observation state in an action chunk. Unlike UMI [15], we use the absolute ee pose $\{\mathbf{p}^{(i)}_{\text{ee}},\mathbf{R}^{(i)}_{\text{ee}}\}$ as the robot state, with the goal of enabling the policy to perceive its pose in the real workspace. Moreover, since our model is designed to provide inference-time steering for both DP [14] and $\pi_{0.5}$ [32], and $\pi_{0.5}$ typically conditions on a single observation step, using relative state would make the state always the identity matrix, which would render proprioception ineffective.
IV Method: TouchGuide
In this section, we introduce TouchGuide, a two-stage cross-policy multimodal fusion paradigm based on classifier guidance [18] (Sec. IV-A), inspired by recent advances in inference-time steering [51, 22, 55]. TouchGuide performs inference-time steering on a pre-trained diffusion-based or flow-matching-based (e.g., action expert of $\pi_{0.5}$) visuomotor policy $\pi_{\theta}(\mathbf{A}_{t}\,|\,\mathbf{V}_{t})$. During the early phase of the denoising or flow-matching process, the policy generates an initial coarse action based solely on visual observation, ensuring visual feasibility without yet enforcing tactile constraints. In the later phase, a task-specific Contact Physical Model (CPM, detailed in Sec. IV-B) provides tactile guidance via a feasibility score, which steers the action toward better satisfaction of real-world physical contact. The complete inference-time guidance process of TouchGuide is detailed in Sec. IV-C.
IV-A Preliminaries: Classifier Guidance
Diffusion Model. Following the notation in prior work [18], classifier guidance can be written as follows:
$\hat{\epsilon}_{\theta}(x_{t})=\epsilon_{\theta}(x_{t})-\eta\sqrt{1-\bar{\alpha}_{t}}\nabla_{x_{t}}\log p_{\phi}(y\mid x_{t})$ (3)
Flow Matching. Since prior work has not explicitly derived classifier guidance for flow matching [77, 23], we provide a brief introduction to the corresponding flow-matching formulation of classifier guidance.
Proposition 1.
The classifier guidance for flow matching can be formulated as follows:
$\hat{u}_{\theta}(x_{t})=u_{\theta}(x_{t})+\eta\frac{t}{1-t}\nabla_{x_{t}}\log p_{\phi}(y\mid x_{t})$ (4)
where $u_{\theta}(x_{t})$ is the unconditional velocity field, $\phi$ parameterizes the auxiliary noise-dependent classifier, $y$ is the target class label (e.g., additional condition), and $\eta$ is the guidance scale. The term t/(1t)t/(1-t) acts as a time-dependent weighting coefficient derived from the specific flow definition. See the detailed deviration in Appendix -A.
Remark 1.
For VLAs, we can achieve tactile modality fusion by steering their action expert. Specifically, since the action expert of $\pi_{0.5}$ is flow-matching based, we can apply Eq. 4 to steer the action sampling process (See Eq. 12 in Sec. IV-C).
IV-B Contact Physical Model
Algorithm 1 TouchGuide (Inference Time)
1:Base policy $\pi_{\theta}$, Contact Physical Model $s_{\phi}$, total steps $K$, guidance steps $K_{\text{TouchGuide}}$, guidance scale $\eta$, visual observation $\mathbf{V}_{t}$, tactile observation $\mathbf{T}_{t}$
[…]
3:for $k=K,\dots,1$ do
4:  if $\pi_{\theta}\in$ Diffusion-based then $\triangleright$ e.g., DP, DP3
5:   $\epsilon\leftarrow\pi_{\theta}(\mathbf{V})$
6:   if $k\leq K_{\text{TouchGuide}}$ then
7:     $\mathbf{s}\leftarrow s_{\phi}(\mathbf{V}_{t},\mathbf{T}_{t},\mathbf{A}^{k}_{t})$
8:     $\hat{\epsilon}\leftarrow\epsilon-\eta\sqrt{1-\bar{\alpha}_{k}}\nabla_{\mathbf{A}^{k}_{t}}\mathbf{s}$
9:     $\mathbf{A}^{k-1}_{t}\leftarrow$ Update $\mathbf{A}^{k}_{t}$ using $\hat{\epsilon}$
10:   else
11:     $\mathbf{A}_{t}^{k-1}\leftarrow$ Update $\mathbf{A}_{t}^{k}$ using $\epsilon$}
12:   end if
13:  else if $\pi_{\theta}\in$ Flow-matching then $\triangleright$ e.g., $\pi_{0.5}$
14:   $k\leftarrow$ Scale $[K,0]$ to $[1,0]$
15:   $u\leftarrow\pi_{\theta}(\mathbf{V}_{t})$
16:   if $k\leq K_{\text{TouchGuide}}$ then
17:     $\mathbf{s}\leftarrow s_{\phi}(\mathbf{V}_{t},\mathbf{T}_{t},\mathbf{A}_{t}^{k})$
18:     $\hat{u}\leftarrow u+\eta\frac{k}{1-k}\nabla_{\mathbf{A}_{t}^{k}}\mathbf{s}$
19:     $\mathbf{A}_{t}^{k-1}\leftarrow$ Update $\mathbf{A}_{t}^{k}$ using $\hat{u}$
20:   else
21:     $\mathbf{A}_{t}^{k-1}\leftarrow$ Update $\mathbf{A}_{t}^{k}$ using u{u}
22:   end if
23:  end if
24:end for
25:return $\mathbf{A}_{t}^{0}$
At time step $t$, Contact Physical Model (CPM, shown in Fig. 3(a)) outputs a task-specific feasibility score conditioned on the tactile feedback $\mathbf{T}_{t}$, the visual observation $\mathbf{V}_{t}$ and the coarse action $\mathbf{A}_{t}$ generated by the early phase of pre-trained policy $\pi_{\theta}(\mathbf{A}_{t}\,|\,\mathbf{V}_{t})$. Specifically, we use pretrained DINOv2 as both the tactile and visual encoder to obtain the tactile and visual embedings, i.e., $\mathbf{T}_{t}^{\text{emb}}$ and $\mathbf{V}^{\text{emb}}_{t}$ respectively, as it provides effective latent-space representations of robotic observations [22, 78]. Then, we fuse these two embedings using N-layer Transformer encoder $\mathtt{TranEnc}(\cdot)_{\times\text{N}}$ and finally, yield the latent observation $\mathbf{O}_{t}$:
$\mathbf{O}_{t}^{\text{emb}}=\mathtt{TranEnc}(\mathtt{concat}[\mathbf{T}_{t}^{\text{emb}},\mathbf{V}_{t}^{\text{emb}}])_{\times\text{N}}\xrightarrow[]{\text{L2 Norm}}\mathbf{O}_{t}$ (5)
Similarly, we can obtain the latent action $\mathbf{a}_{t}$ by:
$\mathbf{A}_{t}^{\text{emb}}\xrightarrow[]{\text{L2 Norm}}\mathbf{a}_{t}$ (6)
where, $\mathbf{A}_{t}^{\text{emb}}$ is the action embedding generated by an action encoder taking $\mathbf{A}_{t}$ as input. The feasibility score $\mathbf{s}$ represents the cosine similarity between the observation and the action in the latent space [11], and is defined as follows:
$\mathbf{s}=s_{\phi}(\mathbf{V}_{t},\mathbf{T}_{t},\mathbf{A}_{t})=\mathbf{O}_{t}^{\top}\mathbf{a}_{t}$ (7)
Training Strategy. We use the limited expert demonstrations to train our task-specific CPM. Given the visual-tactile observation (i.e., $\mathbf{V}_{t}$ and $\mathbf{T}_{t}$) and an ground-truth action $\mathbf{A}_{t}$, we can yield the latent observation $\mathbf{O}_{t}$ and latent action $\mathbf{a}_{t}$. Subsequently, we establish the positive samples using latent observation-action pairs from the same time step $\{(\mathbf{O}_{t},\mathbf{a}_{t})\}$, and negative samples using pairs from different time steps $\{(\mathbf{O}_{t+\Delta t},\mathbf{a}_{t}),(\mathbf{O}_{t},\mathbf{a}_{t+\Delta t})\}$ . Finally, a standard contrastive learning loss $\mathcal{L}_{\text{CPM}}$ is adopted:
$\mathcal{L}_{\text{CPM}}(\mathbf{O}_{t},\mathbf{a}_{t})=\frac{1}{2}(\mathcal{L}_{\mathbf{O}\to\mathbf{a}}+\mathcal{L}_{\mathbf{a}\to\mathbf{O}})$ (8)
$\mathcal{L}_{\mathbf{O}\to\mathbf{a}}$ $=-\frac{1}{M}\sum_{i=1}^{M}\log\frac{\exp(\mathbf{O}_{i}^{\top}\mathbf{a}_{i}/\tau)}{\sum_{j=1}^{M}\exp(\mathbf{O}_{i}^{\top}\mathbf{a}_{j}/\tau)}$ (9)
$\mathcal{L}_{\mathbf{a}\to\mathbf{O}}$ $=-\frac{1}{M}\sum_{i=1}^{M}\log\frac{\exp(\mathbf{a}_{i}^{\top}\mathbf{O}_{i}/\tau)}{\sum_{j=1}^{M}\exp(\mathbf{a}_{i}^{\top}\mathbf{O}_{j}/\tau)}$ (10)
where $\tau$ is a learnable temperature parameter and $M$ denotes the number of samples used during training. In order to enable the CPM to effectively operate in a noisy action space (generated by the early phase of the pretrained policy) during inference, we enhance the ground-truth action $\mathbf{A}_{t}$ by adding noise following a geometric distribution using the same noise scheduler as DP [22]; for flow-matching policies, we instead apply linear interpolation.
IV-C Steering the Policies via Touch Guidance
Figure 4: Comparison of policy distributions in action space.
Fig. 3(b) illustrates the entire TouchGuide pipeline, where touch guidance is used to steer the sampling process of base policy to generate actions that better adhere to the contact physics. Following [22, 51], both diffusion models and flow-matching models can be guided by the gradients $\nabla_{x_{t}}\log p_{\phi}(y\mid x_{t})$ provided by a classifier toward a particular set of samples that maximizes $p(y\mid x)$ (See Eq. 34[18, 77, 23]. In our case, $\log p_{\phi}(y\mid x_{t})$ corresponds to the feasibility score $\mathbf{s}$ produced by the CPM. With this formulation, touch guidance for flow-matching (e.g., action expert of $\pi_{0.5}$) and diffusion policies can be respectively expressed by the following two equations:
$\hat{\epsilon}_{\theta}(\mathbf{A}_{t}^{k},\mathbf{O}_{t})=\epsilon_{\theta}(\mathbf{A}_{t}^{k},\mathbf{V}_{t})-\eta\sqrt{1-\bar{\alpha}_{k}}\nabla_{\mathbf{A}_{t}^{k}}s_{\phi}(\mathbf{O}_{t},\mathbf{A}_{t}^{k})$ (11)
$\hat{u}_{\theta}(\mathbf{A}_{t}^{k},\mathbf{O}_{t})=u_{\theta}(\mathbf{A}_{t}^{k},\mathbf{V}_{t})+\eta\frac{k}{k-1}\nabla_{\mathbf{A}_{t}^{k}}s_{\phi}(\mathbf{O}_{t},\mathbf{A}_{t}^{k})$ (12)
where $\eta$ is the guidance scale, $\epsilon_{\theta}(\cdot)$ and $u_{\theta}(\cdot)$ are the noise predictor in diffusion policy and velocity predictor in flow-matching policy, respectively. The detailed TouchGuide procedure is provided in Alg. 1.
As shown in Fig. 3(c), our TouchGuide can be viewed as using the CPM to steer the policy distribution $\mathcal{Q}_{\text{policy}}$ toward the real distribution $\mathcal{Q}_{\text{real}}$, which refers to the optimal distribution to accomplish the task in the real world. Therefore, sampling actions from the CPM-steered $\mathcal{Q}_{\text{policy}}$ is more likely to produce actions that better adhere to the contact physics. In Fig. 4, we illustrate the policy distributions of TouchGuide and previous methods (including feature-level concatenation [65, 57, 13, 4] and policy-level composition [8, 9]) in the action space. For feature-level concatenation, although tactile sensing provides additional information, the model may still overly attend to the visual modality (since tactile signals often exhibit limited variation in many cases), and thus its policy distribution $\mathcal{Q}_{\text{policy}}$ cannot be explicitly driven toward the real distribution $\mathcal{Q}_{\text{real}}$. Policy-level composition, though an improvement over feature-level methods, still falls short in enabling effective multi-modal fusion within the action space $\mathcal{A}$. Its distribution operates multiplicatively in $\mathcal{A}$, and it lacks an explicit mechanism to steer the $\mathcal{Q}_{\text{policy}}$ toward $\mathcal{Q}_{\text{real}}$. To address these limitations, we introduce a CPM. Our method assesses the physical feasibility of both tactile observations and noisy actions, thereby providing explicit guidance for steering in $\mathcal{A}$.
V Experiments
Figure 5: Five experiment tasks including Shoe Lacing, Chip Handover, Cucumber Peeling, Vase Wiping, and Lock Opening.
In this section, we evaluate the proposed TouchGuide framework and the TacUMI data-collection system across five challenging tasks. Specifically, our experiments are designed to answer the following questions: Q1. How does tactile feedback improve fine-grained robotic manipulation? Q2. How does TouchGuide perform compared to prior visuo-tactile policies? Q3. Does TouchGuide generalize across different robots, tactile sensors, and policies? Q4. Can TacUMI efficiently provide high-quality data for precise contact-rich tasks?
TABLE I: Policy performance comparison on five challenging fine-grained, contact-rich tasks.
Bi-Arx5 (Dual-arm) Flexiv Rizon4 (Single-arm)
Shoe Lacing (100 Demos) Chip Handover (50 Demos) Cucum. Peeling (50 Demos) Vase Wiping (30 Demos) Lock Opening (20 Demos) Average
DP Diffusion Policy 0% 5% 0.500 0.265 0% 16.3%
DP w/ Tactile Observation 0% 15% 0.520 0.240 5% 19.2%
SafeDiff 0% 10% 0.635 0.325 5% 22.2%
RDP 0% 20% 0.740 0.475 10% 30.3%
Policy Consensus 0% 15% 0.630 0.405 5% 24.7%
Tactile Dynamics 0% 15% 0.695 0.335 0% 23.6%
TouchGuide (Force) 0% 30% 0.805 0.510 15% 35.3%
TouchGuide (Tactile Img.) 0% 25% 0.810 0.550 20% 36.2%
$\pi_{0.5}$ $\pi_{0.5}$ 20% 25% 0.785 0.360 20% 35.9%
$\pi_{0.5}$ w/ Tactile Observation 20% 15% 0.760 0.305 10% 30.3%
Tactile Dynamics 10% 30% 0.815 0.395 20% 36.2%
TouchGuide (Force) 25% 40% 0.955 0.590 30% 49.9%
TouchGuide (Tactile Img.) 35% 60% 0.975 0.675 30% 58.0%
V-A Experimental Setup
V-A1 Hardware
For dual-arm tasks, we employ the Bi-ARX5 [2], while the Flexiv Rizon4 [24] integrated with TacUMI is used for single-arm tasks. Given the high symmetry of the gripper’s tactile sensing, we attach tactile sensors only to the outer finger of each arm during dual-arm experiments. This design reduces dataset size and enhances training efficiency.
V-A2 Tasks
As shown in Fig. 5, we design five challenging contact-rich tasks. To thoroughly evaluate the performance of our paradigm, these tasks include long-horizon, precise manipulation (e.g., Shoe Lacing), collaborative handover of fragile objects (e.g., Chip Handover), and low-data (only 20 demos), high-precision manipulation (e.g., Lock Opening). More detailed task descriptions and examples of the evaluation metrics are provided in Appendix -H.
Shoe Lacing. The robot needs to grasp the shoelace and sequentially thread it through four eyelets (two on each side). The task is considered successful if the robot threads the shoelace through all four eyelets within the allotted time.
Chip Handover. The robot needs to grasp a potato chip, execute a collaborative in-air handover, and finally place it into a designated container. The task is considered successful if, within the allotted time, the chip is placed in the container without any breaking.
Cucumber Peeling. The robot must grasp both a peeler and a cucumber, after which the two robotic arms collaborate to peel the cucumber. Task score is assessed based on the ratio of the length of peel removed by the robot to that removed by a human expert. The resulting score is quantized into one of seven discrete values: 0, 0.1, 0.3, 0.5, 0.7, 0.9, or 1.
Vase Wiping. The robot needs to pick up a blackboard eraser from the table, approach the vase held by a human, and erase the marker traces on its surface. The task score is defined similarly to Cucumber Peeling.
Lock Opening. The robot executes a sequential manipulation task involving a human partner: grasping a key held by the human, inserting it into a lock also held by the human, and finally turning the key to open the lock. The task is considered successful if the robot maintains hold of the key without dropping it and opens the lock within the allotted time.
V-A3 Baselines
We adopt two state-of-the-art (SOTA) visuomotor policies (i.e., DP [14] and $\pi_{0.5}$ [32]) as the base policy for TouchGuide, and use several visuo-tactile policies (e.g., RDP [65], Policy Consensus [9] and SafeDiff [57]) as the baselines for comparison. Specifically, we include Tactile Dynamics as one of our baselines. This method is derived from Latent Policy Barrier [51] and employs a dynamics model to jointly predict tactile and visual observations. Details on the touch guidance setup of Tactile Dynamics for $\pi_{0.5}$ can be found in Sec. IV-A. For TouchGuide, we use two tactile modalities (i.e., force and tactile image) to thoroughly validate the cross-modality capability of our paradigm. The specific baseline configurations are detailed in Appendix -G. For these five tasks, each baseline is evaluated over 20 trials with moderate randomization of the initial environment conditions, and the results are reported in Table I.
V-B Main Results and Analysis
Tactile sensing provides rich and accurate in-hand state information, which is crucial for fine-grained manipulation (Q1). Tactile sensing not only provides rich force information but also yields accurate in-hand state information. For fine-grained manipulation tasks such as shoe lacing, chip handover, and lock opening, this in-hand state information is particularly critical, as it determines the end-effector pose required for threading through an eyelet or executing a reliable handover. In contrast, such in-hand states are difficult to infer accurately from vision alone and are often only weakly observable through specular highlights. Table I shows that TouchGuide yields clear performance gains over DP and $\pi_{0.5}$ (16.3% $\rightarrow$ 36.2%, 35.9% $\rightarrow$ 58.0%). This highlights the importance of accurate in-hand state estimation for precise manipulation.
TouchGuide can explicitly steers base policy in the action space using touch guidance, and it achieves clear performance gains over current SOTA visuo-tactile policies (Q2). Existing methods primarily focus on reacting promptly to changes in tactile signals (e.g., RDP [65] and PolicyConsensus [9]). However, in many tasks, the initial contact often plays a decisive role in task success. For tasks such as lock opening and shoe lacing, small errors at first contact can cause the key to drop or become severely misaligned, driving the system into states that are entirely unfamiliar to the policy, which is particularly detrimental for imitation learning. In contrast, our method leverages a Contact Physical Model to assess the feasibility of the noisy actions proposed by the base policy and uses the resulting feasibility scores to steer the policy at inference time. This paradigm of evaluating the physical consistency between noisy actions and tactile signals is crucial for producing appropriate contacts. Table I supports this claim, showing that TouchGuide (36.2%) achieves a noticeable performance improvement over other SOTA visuo-tactile policies (e.g., RDP [65] (30.3%) and PolicyConsensus [9] (24.7%)).
TouchGuide demonstrates strong generalization, and can be applied across different robots, tactile modalities, and policies (Q3). Unlike prior visuo-tactile policies, TouchGuide augments a visuomotor policy with touch guidance in a fully decoupled manner, requiring no retraining to the base policy. As a result, it can inherit the strong capability of the base policy. For example, TouchGuide built on $\pi_{0.5}$ achieves solid performance on a high-precision task such as lock opening with only 20 demonstrations. Moreover, TouchGuide achieves strong performance across different robots (e.g., Bi-ARX5 and Flexiv Rizon4), different policies (e.g., DP and $\pi_{0.5}$), and different tactile modalities (e.g., force and tactile image). Detailed results are reported in Table I. More baseline failure cases are provided in Appendix -C.
V-C Ablation Studies
V-C1 Noise Pretraining Ablation
Since the task-specific CPM takes noisy actions as input at inference time, we add noise into the action using the same noise scheduler for diffusion models, or via linear interpolation for flow-matching models. The number of noising steps is sampled from a geometric distribution. As shown in Table II, this key setting can effectively improve the task success rate for TouchGuide (39.17% $\rightarrow$ 62.50%).
TABLE II: Ablation study on noise pretraining.
Chip Cucumber Lock
Handover Peeling Opening Average
Ours w/o Noise Pretraining 30% 0.725 15% 39.17%
Ours w/ Noise Pretraining 60% 0.975 30% 62.50%
V-C2 CPM Modality Ablation
For most fine-grained manipulation tasks, tactile and visual signals are sparse at different time periods. For example, in the Cucumber Peeling task, tactile feedback is sparse when the robot holds the peeler and moves toward the cucumber. During peeling, however, small motion variations lead to pronounced changes in tactile feedback, while visual information becomes relatively sparse. Therefore, we argue that for task-specific CPM, the visual and tactile modalities are equally important (see Table III). The results show without vision, the success rate drops from 62.50% to 43.33%, and without touch, it decreases to 43.50%.
TABLE III: Ablation study on CPM input modalities.
Chip Cucumber Lock
Handover Peeling Opening Average
Ours w/o Vision 30% 0.850 15% 43.33%
Ours w/o Touch 35% 0.755 20% 43.50%
Ours w/ Vision & Touch 60% 0.975 30% 62.50%
Also,we provide ablation studies on the guidance scale $\eta$ and guidance steps $K_{\text{TouchGuide}}$ in Alg. 1 (See Appendix -B).
V-D System Comparison and User Study
TABLE IV: Success rate comparison of data collection systems on the Lock Opening task.
Method TacUMI UMI (SLAM-based) VR Teleop
$\pi_{0.5}$ 20% 0% 5%
TouchGuide 30% 5% 15%
TacUMI provides accurate and reliable data that is well suited for robot learning (Q4). We find that although VR teleoperation provides highly precise action (since it is read directly from the robot arm), teleoperating tasks, such as Lock Opening, are challenging for human demonstrators. Hesitation around the keyhole often degrades the data quality, making it unsuitable for robot learning. This issue is particularly pronounced in the low-data regime (20 demos), where such low-quality demonstrations can even hinder learning. In addition, SLAM-based UMI [15] struggles to meet the requirements of fine-grained manipulation tasks due to the limited accuracy of its SLAM module. In order to evaluate the effectiveness of our TacUMI, we trained policies using the data collected by all involved data collection systems and evaluated them on the Lock Opening task. Table IV indicates that our TacUMI outperforms the other two data collection systems.
To evaluate the usability of our TacUMI, we recruited five volunteers working in robotics (two had prior experience collecting demonstrations using a handheld gripper, while three had no such experience). Each volunteer collected 20 valid demonstrations using VR Teleop and UMI with three different tracking methods (e.g., SLAM-based, AR motion tracker, and TacUMI). We used the following metrics: Attempts (the number of trials required to obtain the target number of demonstrations), Length (the dataset duration excluding preparation and post-processing time, which helps assess system usability), Valid Rate (Valid Data / Attempts), and Satisfaction (each volunteer’s subjective usability rating of the system). As shown in Table V, although Teleop achieves a high valid rate, the tactile feedback is not directly available, which leads to excessively long demonstrations and may degrade policy performance. SLAM-based tracking yields a very low valid rate. The AR Motion Tracker requires wearing an AR headset, making it unsuitable for prolonged data collection, and occlusions can easily cause tracking loss if the user is not constantly attentive during collection. TacUMI demonstrates excellent usability and receives high satisfaction ratings from the volunteers. For a comparative analysis with alternative data collection systems, please refer to Appendix -D1.
TABLE V: User study on data collection systems.
Method Attempts $\downarrow$ Length $\downarrow$ Valid Rate $\uparrow$ Satisfaction $\uparrow$
VR Teleop 100 23.0 min 100.00% 7.8
SLAM-based 143 15.3 min 69.93% 7.2
AR Motion Tracker 104 14.7 min 96.15% 9.0
TacUMI (Ours) 100 13.5 min 100.00% 9.6
VI Limitations and Future Work
Although our TouchGuide paradigm enables tactile fusion without retraining the base policy, the CPM still requires task-specific training, which inevitably increases training time and cost. An important direction for future work is to learn more effective tactile representations and develop a non-task-specific CPM that can generalize across tasks without retraining.
VII Conclusion
In this paper, we propose a novel two-stage cross-policy multi-modal fusion paradigm (i.e., TouchGuide), and an efficient data collection system (i.e., TacUMI). TouchGuide steers the base policy via touch guidance to generate actions that better adhere to the contact physics, demonstrating strong performance on five challenging fine-grained manipulation tasks. TacUMI strikes an trade-off between cost and precision, minimizing overhead while maintaining high accuracy; moreover, its direct tactile feedback improves usability and helps operators collect high-quality tactile data more efficiently.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (NSFC) under Grants No. 12522202, No. 12272220 and No. 124B1038, and in part by Xense Robotics.